### v1

# Is Your Paper Being Reviewed by an LLM? A New Benchmark Dataset and Approach for Detecting AI Text in Peer Review

Abstract
Peer review is a critical process for ensuring the integrity of published scientific research. Confidence in this process is predicated on the assumption that experts in the relevant domain give careful consideration to the merits of manuscripts which are submitted for publication. With the recent rapid advancements in large language models (LLMs), a new risk to the peer review process is that negligent reviewers will rely on LLMs to perform the often time consuming process of reviewing a paper. However, there is a lack of existing resources for benchmarking the detectability of AI text in the domain of peer review. To address this deficiency, we introduce a comprehensive dataset containing a total of 788,984 AI-written peer reviews paired with corresponding human reviews, covering 8 years of papers submitted to each of two leading AI research conferences (ICLR and NeurIPS). We use this new resource to evaluate the ability of 18 existing AI text detection algorithms to distinguish between peer reviews written by humans and different state-of-the-art LLMs. Motivated by the shortcomings of existing methods, we propose a new detection approach which surpasses existing methods in the identification of AI written peer reviews. Our work reveals the difficulty of identifying AI-generated text at the individual peer review level, highlighting the urgent need for new tools and methods to detect this unethical use of generative AI.
Is Your Paper Being Reviewed by an LLM? A New Benchmark Dataset and Approach for Detecting AI Text in Peer Review
Sungduk Yu Intel Labs sungduk.yu@intel.com                        Man Luo Intel Labs man.luo@intel.com                        Avinash Madusu Intel Labs avinash.madasu@intel.com
Vasudev Lal Intel Labs vasudev.lal@intel.com                        Phillip Howard Intel Labs phillip.r.howard@intel.com
1 Introduction
Figure 1: Left panel shows the data construction pipeline and the right panel shows the proposed method Anchor Embedding Approach for detecting AI generated review.
Recent advancements in large language models (LLMs) have enabled their application to a broad range of domains, where LLMs have demonstrated the ability to produce plausible and authoritative responses to queries even in highly technical subject areas. These advancements have coincided with a surge in interest in AI research, resulting in increased paper submissions to leading AI conferences (Audibert et al., 2022). Consequently, workloads for peer reviewers have also increased significantly, which could make LLMs an appealing tool for lessening the burden of fulfilling their peer review obligations (Kuznetsov et al., 2024; Kousha and Thelwall, 2024; Zhuang et al., 2025).
Despite their impressive capabilities, the use of LLMs in the peer review process raises several ethical and methodological concerns which could compromise the integrity of the publication process (Hosseini and Horbach, 2023; Latona et al., 2024; Seghier, 2024; Zhou et al., 2024). Reviewers are selected based on their expertise in a technical domain related to a submitted manuscript, which is necessary to critically evaluate the proposed research. Offloading this responsibility to an LLM circumvents the role that reviewer selection plays in ensuring proper vetting of a manuscript. Furthermore, LLMs are prone to hallucination and may not possess the ability to rigorously evaluate research publications. Therefore, the use of LLMs in an undisclosed manner in peer review poses a significant ethical concern that could undermine confidence in this important process.
Motivating the need for evaluation resources and detection tools to address this problem is the apparent increase in AI-generated text among peer reviews submitted to recent AI research conferences. Recent studies revealed an upward trend in AI-generated texts among peer reviews at the corpus level (Liang et al., 2024a; Latona et al., 2024). This trend is particularly concerning given that evaluations from human and AI reviewers are not aligned (Drori and Te’eni, 2024; Latona et al., 2024; Ye et al., 2024), suggesting that the unregulated and undisclosed use of LLMs in peer review could undermine the integrity of the current system.
Despite the growing recognition of this problem, there is a lack of existing dataset resources for comprehensively evaluating the performance of AI text detection methods in the domain of peer review. To address this deficiency, we introduce the largest dataset to-date of parallel human-written and LLM-written peer reviews for identical papers submitted to 8 years of papers for two leading AI research conferences, NeurIPS and ICLR (Figure 1). Our dataset consolidates human-written peer reviews from existing sources with AI-written peer reviews that we generated for the same paper using five state-of-the-art LLMs: GPT-4o Achiam et al. (2023), Claude Sonnet 3.5 Anthropic (2023), Gemini 1.5 pro Team et al. (2023), Qwen 2.5 72b Bai et al. (2023), and Llama 3.1 70b Dubey et al. (2024). In total, our dataset contains 788,984 peer reviews, evenly balanced between human-written reviews and AI-generated peer reviews created by these five LLMs.
We use our dataset to investigate the suitability of various AI text detection methods for identifying LLM generations in the peer review process. While limited prior work has analyzed the presence of AI-generated text in peer reviews at the corpus level (Liang et al., 2024a) or used conference-submitted reviews only (Latona et al., 2024), our study is the first to investigate the detectability LLM generations at the individual review level using systematically generated AI samples along with human samples, which is necessary to address this problem in practice. Specifically, we evaluate 18 open-source methods.
Our results show that most existing AI-text detection methods are limited in their ability to robustly detect AI-generated reviews while maintaining a low number of false positives. Motivated by this finding, we propose an alternative approach to detecting AI-generated peer reviews by comparing the semantic similarity of a given review to a set of reference AI-generated reviews for the same paper, which surpasses the performance of all existing approaches in detecting GPT-4o and Claude written peer reviews. We conduct analyses to understand how different levels of AI use for editing reviews impacts detectability and false positives, as well as the characteristics which distinguish LLM-written peer reviews from those written by humans. Our work demonstrate the challenge of detecting AI-written text in peer reviews and motivates the need for further research on methods to address this unethical use of LLMs in the peer review process.
To summarize, our contributions are as follows: (1) We introduce a new dataset containing 788,984 AI-written peer reviews generated by five widely-used LLMs along with corresponding human-written peer reviews for the same papers, which is the largest resource of its kind for evaluating the detectability of AI text in peer review. (2) Using our dataset, we evaluate 18 existing open source AI text detection algorithms as well as a commercial API, finding that most methods struggle to reliably detect fully AI-written peer reviews at low levels of false positives. (3) We propose a new detection method which compares the semantic similarity of a given peer review to a reference LLM-generated peer review for the same paper, which outperforms all existing detection approaches in our evaluations. (4) We conduct analyses to highlight the characteristics which distinguish human-written peer reviews from those generated by LLMs, show LLM reviews are generally more favorable and more confident it their assessment, and evaluate how the use of LLMs for editing peer reviews impact false positive rates. (5) We make our dataset publicly available under an open source license to facilitate future work on detecting AI generated text in peer review.
2 Related Work
AI text detection datasets
Several benchmark datasets have been introduced to evaluate AI-based text detection models. RAID-TD  Dugan et al. (2024) provides a large-scale benchmark designed to assess text detection under adversarial conditions, ensuring robustness against manipulated AI-generated content. The M4 Dataset  Wang et al. (2024) expands the scope by incorporating reviews from multiple LLMs across different languages, offering a more diverse linguistic evaluation. The HC3 Dataset  Guo et al. (2023) consists of responses from ChatGPT and human experts, covering specialized domains such as finance, medicine, and law, in addition to general open-domain content. In contrast, the GPT Reddit Dataset (GRiD)  Qazi et al. (2024) focuses on social media conversations, compiling a diverse set of human- and AI-generated responses to Reddit discussions. Meanwhile, Beemo Artemova et al. (2024) introduces a benchmark of expert-edited machine-generated outputs, spanning creative writing, summarization, and other practical applications, further refining detection challenges. These benchmarks primarily evaluate AI-generated text from a single model and do not address the domain of AI text in peer review. In contrast, our dataset is larger than most existing datasets (788k generations) and is unique in its focus on AI text detection in peer review.
AI-generated text detection
AI-generated text detection has been framed as a binary classification task to distinguish human-written from machine-generated text  Bakhtin et al. (2019); Jawahar et al. (2020); Fagni et al. (2021); Mitchell et al. (2023a). Solaiman et al. (2019) used a bag-of-words model with logistic regression for GPT-2 detection, while fine-tuned language models like RoBERTa Liu et al. (1907) improved accuracy  Zellers et al. (2019); Uchendu et al. (2020); Gehrmann et al. (2019). Zero-shot methods based on perplexity and entropy emerged as alternatives  Ippolito et al. (2020); Gehrmann et al. (2019). Other studies focused on linguistic patterns and syntactic features for model-agnostic detection  Uchendu et al. (2020); Gehrmann et al. (2019). Watermarking techniques, such as DetectGPT  Mitchell et al. (2023a), have also been proposed for proactive identification. Centralized frameworks like MGTBench  He et al. (2023) and its refined version, IMGTB  Spiegel and Macko (2023), provide standardized evaluations for AI text detection. IMGTB categorizes methods into model-based and metric-based approaches. Model-based methods leverage large language models such as ChatGPT-turbo  OpenAI (2023) and Claude  Anthropic (2023). Metric-based methods, including Log-Likelihood  Solaiman et al. (2019), Rank  Gehrmann et al. (2019), Entropy  Gehrmann et al. (2019), DetectGPT  Mitchell et al. (2023a), and LRR  Su et al. , rely on log-likelihood and ranking for classification.
AI-assisted peer review
Recent studies have explored the role of LLMs in peer review, examining their influence on reviewing practices  Liang et al. (2024a, b), simulating multi-turn interactions  Tan et al. (2024), and assessing their reviewing capabilities  Zhou et al. (2024). Tools like OpenReviewer  Tyser et al. provide AI-assisted review improvements, while other works focus on LLM transparency  Kuznetsov et al. (2024) and distinguishing AI-generated content  Mosca et al. (2023). Other recent work has investigated AI-driven review systems  Tyser et al. (2024), agentic frameworks  Liang et al. (2024b), and comparative analyses of LLM-generated reviews  Liu and Shah (2023), along with broader explorations of LLMs’ role in peer review  Jin et al. (2024); Santu et al. (2024); Sukpanichnant et al. (2024). While it is not the primary focus of our work, we analyze the quality of LLM-generated peer reviews in Section 6.
Calibration set Test set Extended set
ICLR NeurIPS ICLR NeurIPS ICLR NeurIPS
GPT-4o 7,710 7,484 27,342 30,212 121,278 91,994
Gemini 1.5 Pro 7,704 7,476 27,278 30,146 - -
Claude Sonnet 3.5 v2 7,672 7,484 27,340 30,208 - -
Llama 3.1 70b 7,708 7,472 27,284 30,086 121,130 91,706
Qwen 2.5 72b 7,662 7,452 27,190 29,966 - -
Total 75,824 287,052 426,108
Grand total 788,984
Table 1: Dataset statistics. Each subset in the dataset is balanced, containing an equal number of human-written and AI-generated peer reviews. The calibration and test sets include reviews generated by five different large language models (LLMs), whereas the extended set includes reviews generated by GPT-4o and Llama 3.1.
3 Methods and Dataset Construction
3.1 Human reviews
We used the OpenReview API (OpenReview, ) to collect submitted manuscripts and their reviews for the ICLR conferences from 2019 to 2024, as well as for NeurIPS conferences from 2021 to 2024. Additionally, we used the ASAP dataset Yuan et al. (2022) to collect manuscripts and reviews for ICLR 2017 to 2018 and NeurIPS 2016 to 2019111NeurIPS 2020 reviews are not publicly available.
3.2 AI reviews
We generated 788,984 AI-generated reviews using five widely-used LLMs: GPT-4o, Claude Sonnet 3.5, Gemini 1.5 pro, Qwen 2.5 72b, and Llama 3.1 70b.
Prompts. To control the content and structure of these AI-generated reviews, we included conference-specific reviewer guidelines and review templates in the prompts. Note that this is nontrivial since the review templates have evolved significantly over the years (Table 5). In addition, we aligned the paper decisions by prompting the LLMs with specific decisions derived from the corresponding human reviews (see Appendix E for complete prompt details). This step is important, as we found that AI review content and recommendations vary substantially depending on how input prompts are constructed. Thus, these measures represent our efforts to control the influence of text prompts.
Computation. We used Azure OpenAI Service, Amazon Web Services, and Google Cloud Platform to generate GPT-4o, Claude, and Gemini reviews, respectively. For Qwen reviews, we used NVIDIA RTX 6000 GPUs, and for Llama reviews, we used Intel Gaudi-2 accelerators.
3.3 Dataset Statistics
Table 1 provides complete statistics for our generated dataset (see Appendix B.3 for a breakdown by conference year and review-generating LLM). We withheld a randomly sampled subset of reviews to serve as a calibration set, which is used in our experiments to determine classification thresholds for each evaluated method. This calibration set contains 75,824 AI-generated and human-generated peer reviews, divided approximately evenly across all five LLMs. To construct the calibration set, we randomly selected 500 papers from ICLR (2021, 2022) and NeurIPS (2021, 2022) and generated AI reviews corresponding to the human reviews for each paper. Because our sampling was done at the paper level rather than the review level, the number of reviews per paper—and consequently per conference—varies slightly. To facilitate the evaluation of detection methods which are more computationally expensive (e.g., methods which requires using LLMs as surrogates), we also withheld a separate test set consisting of human reviews and those generated by all five LLMs for 500 randomly sampled papers from each conference & year. This test split contains a total of 287,052 reviews and is used throughout our main experimental results. A further 426,108 reviews were generated from GPT-4o and Llama 3.1 70b, which we designate as an extended set of reviews (additional experimental results for the extended set are provided in Appendix C.3).
We note slight variations in sample sizes among the five LLMs. These differences stem from LLM-specific factors such as context window limits, generation errors (e.g., malformed JSON or gibberish outputs), and input prompt safety filters.
3.4 Review Editing
LLMs have been widely used for editing human writing Laban et al. (2024); Raheja et al. (2024), including tasks such as grammar checking and enhancing language fluency, which are particularly beneficial for non-native speakers in conveying their reviews more clearly. However, some users may overly rely on LLMs, leading to significant alterations of the original review. To simulate such scenarios in our dataset, we generate edited reviews with four varying levels of modification using LLMs (minimum, moderate, extensive, and maximum). This approach enables us to assess how robust detection methods are for text with different degrees of editing. Appendix E.3 illustrates the types of prompts used to generate these edits.
To validate whether these prompts produce reviews with varying levels of modification, we conducted a similarity checking experiment. Our hypothesis is that the similarity between the edited review and the original review decreases as the level of modification increases, ranging from minimal edits to extensive revisions. To test this, we used a sentence embedding model to compute embeddings for both the original and edited reviews and calculated their average cosine similarity scores. The results, summarized in Table 2, confirm that the editing levels are indeed distinct.
Minimum Moderate Extensive Maximum
0.9841 0.9261 0.8616 0.6799
Table 2: Similarity score of edited reviews and original human-written reviews using different prompts.
4 Anchor Embeddings Detection Methodology
We propose a novel method for detecting AI-generated reviews by comparing their semantic similarity to a reference AI-generated review for the same article, referred to as the “Anchor Review" (AR). The AR can be generated by any LLM. We use a simple, generic prompt (Appendix E.2) to generate the AR without prior knowledge of the user prompts (i.e., the AR prompt differs from those used to create reviews in the testing dataset). Once an AR is generated for a given paper (Eq.1), and a testing review (TR) is provided, we obtain their embeddings using a text embedding model (EM, Eqs.2 and 3). The semantic similarity between the embeddings of the AR and TR is then computed using the cosine similarity function (Eq.4). Finally, this similarity score is compared against a learned threshold ($\theta$): if the score exceeds the threshold, the review is classified as AI-generated; otherwise, it is not (Eq. 5). Putting everything together, the method is formalized as:
$\text{AR}=\text{LLM}(\text{paper},\text{Prompt}_{\text{AR}})$ (1)
[…]
$\text{Label}=\begin{cases}1&\text{if }\text{Score}>\theta,\\ 0&\text{otherwise}.\end{cases}$ (5)
In our study, we use OpenAI’s embedding model (text-embedding-003-small). The threshold $\theta$ is learned from the calibration set. Specifically, for each review in the calibration data, we apply the steps outlined in Eqs. 1 to 4.
Voting of Multiple Anchors
Intuitively, our approach performs best when the anchor embeddings are generated using the same model that produced the test review (source LLM). However, in real-world scenarios, the source LLM is typically unknown (a situation commonly referred to as “black-box” detection scenario). To address this challenge, we propose a voting-based technique. Specifically, we generate multiple anchor embeddings using different types of LLMs (anchor LLMs). For each anchor embedding, we compute the Score (Eq.4) and derive the corresponding label assignment (Eq.5). If at least one anchor embedding assigns a positive label, the final label is positive. Otherwise, the final label is negative.
In our experiments in the following sections, we use three anchor reviews for voting—each generated by GPT-4o, Gemini, and Claude, respectively.
5 Experimental Results
5.1 Fully AI-Written Review Detectability
Figure 2: ROC plots computed from the combined GPT-4o, Gemini, and Claude review calibration dataset, showing results for ICLR (left) and NeurIPS (right); AUC values are shown in parentheses.
We compare our approach with 18 baseline methods from IGMBT (see Appendix A for details) and utilize the calibration set to determine appropriate thresholds for the test data. The threshold is then determined by setting an target False Positive Rate (FPR), which is achieved by adjusting the threshold until the FPR equals the target value. In our setup, we have studied varying thresholds (e.g., 0.1%, 0.5%, and 1%). We focus on low FPR targets because false positive classifications—where human-written reviews are mistakenly identified as AI-generated—carry high stakes, potentially damaging an individual’s reputation. Besides, we focus on AI review text samples generated by three commercial LLMs (GPT-4o, Gemini, and Claude) because these models are more advanced, making the text detection task harder. Also, general users are more likely to choose them over open-source LLMs due to their convenient user interfaces and limited access to the compute resources required for running open-source models.
AI text detection models can be calibrated for varying levels of sensitivity in order to balance the trade-off between true positive and false positive detections. Receiver operating characteristic (ROC) curves are therefore commonly used to compare different methods, as they provide a visualization of the true-positive rate (TPR) which can be achieved by a model when its decision boundary is calibrated for a range of different false positive rates (FPR). Figure 2 provides the ROC curves for our anchor embedding approach and baseline methods, calculated using our GPT-4o, Gemini, and Claude review calibration subset separately for reviews submitted to ICLR2021-2022 (left) and NeurIPS2021-2022 (right). The area under the curve (AUC) is provided for each method in the legend; higher values indicate better detection accuracy across the entire range of FPR values. These plots show that our anchor embedding approach consistently achieves the highest AUC for AI peer review detection among evaluated methods, followed by the baseline method Binoculars.
While ROC curves are useful for comparing the overall performance of different classifiers, only the far left portion of these plots are typically relevant for practical applications of AI text detection models. This is particularly true in the domain of peer review, where the cost of a false positive is high. Reviewers volunteer their time and expertise to sustain this important process; false accusations have the potential to further reduce the availability of reviewers due to disengagement and can also lead to significant reputational harm. Therefore, it is vital that AI text detection systems for peer review be calibrated for a low FPR in order to avoid such negative outcomes.
Prior work has shown that AUC is not necessarily reflective of how models perform at very low FPR values (Yang et al., 2023; Krishna et al., 2024; Tufts et al., ). Therefore, we also report the actual TPR and FPR achieved by different detection methods at discrete low values of target FPR (0.1%, 0.5%, and 1%), which we believe to be of greatest interest for practical applications. The target FPR is used to calculate each method’s classification threshold using our calibration dataset, with the actual TPR and FPR computed over the withheld test dataset. To simulate a more challenging evaluation setting where some of the test reviews are “out-of-domain” in the sense that they come from a different conference than the calibration dataset, we use only the ICLR reviews to calibrate each method (see Section C.1 for in-domain evaluations). Table 3 provides these results separately for the detection of GPT-4o, Gemini, and Claude reviews (Llama and Qwen results are provided in Appendix C.2). Other baseline methods that failed achieve a TPR of at least 1% at a target FPR of 1% (i.e., TPR@1%FPR < 1%) are omitted.
The results in Table 3 show that our anchor embedding approach achieves the highest TPR at low FPR values among evaluated methods for GPT-4o and Claude-generated reviews. This is notable, as GPT-4o and Claude are among the most advanced and widely-used LLMs currently available. The performance difference between our approach and the next best-performing method (Binoculars) is particularly large for GPT-4o reviews and low target FPR settings (e.g., 0.1%), with our anchor embedding approach achieving absolute improvements of up to 50.1% in TPR. Furthermore, our anchor embedding approach exactly achieves the target FPR on the test set for these reviews, whereas other methods exceed the target FPR by as much as 0.7%. Gemini-generated reviews appear to be easier for baseline methods to detect than GPT-4o and Claude reviews, and we observe that our anchor embedding approach performs similarly on these reviews as Binoculars. This suggests that the anchor embedding method provides the greatest benefits over existing methods for reviews which are the most challenging to detect; in fact, our approach performs best for detecting GPT-4o reviews, whereas other strong baselines such as Binoculars perform the worst on these reviews.
Target FPR: 0.1% 0.5% 1%
FPR TPR FPR TPR FPR TPR
GPT-4o Reviews Anchor 0.1 63.5 0.5 83.7 1.0 88.8
Binoculars 0.2 17.1 0.6 33.6 1.0 45.2
MAGE-D 0.1 2.3 0.6 8.8 1.3 14.7
s5 0.1 0.1 0.9 7.2 1.7 17.5
MFD 0.2 0.1 0.8 6.0 1.6 15.6
GLTR 0.1 0.1 0.4 1.9 1.1 5.7
DetectGPT 0.1 0.1 0.6 1.1 1.2 2.3
Gemini Reviews Anchor 0.2 59.7 0.8 80.3 1.3 86.5
Binoculars 0.2 61.5 0.6 78.0 1.0 85.5
s5 0.0 0.2 0.5 9.6 1.1 19.4
MFD 0.1 0.5 0.4 8.9 1.1 18.8
FastDetectGPT 0.1 1.1 0.5 5.8 1.1 10.3
GLTR 0.2 0.5 0.8 5.2 1.8 12.4
DetectGPT 0.1 0.4 0.5 3.5 1.2 7.0
MAGE-D 0.1 0.4 0.6 3.3 1.3 7.0
Loglikelihood 0.0 0.0 0.3 0.1 0.5 1.0
Claude Reviews Anchor 0.1 59.6 0.5 75.8 1.0 81.8
Binoculars 0.2 43.5 0.6 65.8 1.0 77.0
s5 0.0 0.1 0.2 7.6 0.5 17.5
MFD 0.0 0.2 0.1 6.8 0.4 16.5
DetectGPT 0.1 0.5 0.6 5.3 1.2 11.1
GLTR 0.0 0.0 0.2 0.5 0.6 1.8
Table 3: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR reviews from our calibration set at different target FPRs. Highest TPRs are in bold.
Reviews generated by open-source LLMs (see Appendix C.2) show somewhat different trends. For both Llama and Qwen reviews, the Binoculars method achieves near-perfect detection performance. While the Anchor method ranks second for Qwen reviews with slightly lower FPRs, it is not as performant on Llama reviews. This is surprising, as Llama reviews appear easier to detect—e.g., nearly twice as many methods achieve TPR@1%FPR > 10% for Llama reviews compared to Qwen, GPT-4o, and Claude reviews. One possibility is that the semantic similarity between the three anchor reviews generated by higher-quality LLMs (GPT-4o, Gemini, and Claude) and Llama reviews is low enough to overlap with that between the anchor reviews and human reviews, potentially blurring the decision boundary. Although the underlying causes of these differences need further investigation, these findings are less central to our study since most LLM users are likely to rely on commercial models due to their ease of use and superior quality.
Method NDCG Minimum Moderate Extensive Maximum Full
Binoculars 0.94 0.14 0.25 0.34 0.56 0.79
Anchor 0.98 0.05 0.09 0.16 0.87 0.94
Table 4: NDCG and the proportion of reviews flagged as AI-written for different levels of editing.
5.2 Detecting AI-Edited Peer Reviews
We also evaluated the ability of our method to correctly rank reviews in ascending order of their likelihood of being AI-generated. The ground truth order is: human-written, minimally-edited, moderately-edited, extensively-edited, maximally-edited, and fully AI-generated reviews. To measure this, we compute the probability that each review type is AI-generated and rank them accordingly. The ranking is compared with the ground truth using the NDCG metric Järvelin and Kekäläinen (2002), with higher values indicating that the predicted ranking is closer to the ground truth ranking. In addition, we analyze the detection rates for each type of review. For each category, we calculate the percentage of reviews identified as AI-generated. Ideally, the detection rate should be lowest for minimally edited reviews and should progressively increase with the extent of AI involvement.
Table 4 compares our method with the best baseline method from our prior experiments (Binoculars). Our anchor embedding method achieves a higher NDCG score compared to the Binoculars approach, indicating that our method ranks each type of review more accurately and closer to the ground truth. Both methods detect increasing percentages of AI-generated text as the extent of editing increases. This suggests that reviews of more extensive editing are more likely to be detected as AI-generated. For Minimum, Moderate, and Extensive types of edited reviews, our method shows a lower AI-generated detection percentage compared to Binoculars. On the other hand, for Maximum and fully AI-generated reviews, our method shows a higher detection rate than Binoculars. This suggests that our approach is better calibrated to detect reviews where AI involvement is predominant.
6 Analysis
6.1 Human Analysis of Differences Between Human and AI-Written Peer Reviews
To better understand the characteristics which differentiate peer reviews written by humans and LLMs, we conducted a quantitative analysis of 32 reviews authored by humans and GPT-4o for 5 papers submitted to ICLR 2021. Specifically, we read an equal number of human and GPT-4 written reviews for each paper and noted differences in the content between them. A distinguishing characteristic of the analyzed human reviews was that they usually contained details or references to specific sections, tables, figures, or results in the paper. In contrast, peer reviews authored by GPT-4o lacked such specific details, instead focusing on higher-level comments. Another key difference identified in our qualitative analysis was the lack of any specific references to prior or related work in peer reviews generated by GPT-4o. Human authored peer reviews often point out missing references, challenge the novelty of the paper by referencing related work, or suggest specific baselines with references that should be included in the study. In contrast, none of the analyzed GPT-4o reviews contained such specific references to related work. Finally, we found that the vast majority of GPT-4o reviews mentioned highly similar generic criticisms which were not found in human-authored reviews for the same paper. Examples of these issues are provided in Table 13 of Appendix D.1.
Prior work has shown that peer reviews written by GPT-4 and humans have a level of semantic similarity which is comparable to that between different human-authored peer reviews, which has been used to advocate for the usefulness of feedback from GPT-4 in the paper writing process (Liang et al., 2024b). In our qualitative analysis, we found that GPT-4 does indeed generate similar higher-level comments as human reviewers, which could account for this semantic similarity. Despite being generic in nature, we would agree that such feedback could be useful to authors seeking to improve their manuscripts. Nevertheless, we believe that the lack of specificity, detail, and consideration of related work in peer reviews authored by GPT-4 demonstrates that it is not suitable for replacing human domain experts in the peer review process.
6.2 Misalignment Between Human and AI Reviews
In addition to qualitative differences in the content of human and AI-written reviews, we also observe a divergence in numeric scores assigned as part of the review. Figure 3 (Appendix D.2) provides histograms depicting the distribution of score differences for soundness, presentation, contribution, and confidence, which are computed by subtracting scores assigned for each category by human reviewers from those assigned by AI reviewers. AI-written peer reviews were matched with their corresponding human review (aligned by paper ID and overall recommendation) to compute the score differences. Confidence scores range from 1 to 5, while all other categories of scores range from 1 to 4. We focus on reviews from NeurIPS 2022, which were produced prior to the release of ChatGPT. This provides greater confidence that the human-labeled reviews were indeed written by humans, with little to no potential AI influence.
All LLMs produce higher scores than human reviews with a high degree of statistical significance, assessed using a two-sided Wilcoxon signed-rank test (see legend for $p$-values). While the difference between human and AI confidence scores are relatively consistent across all three LLMs, Claude exceeds human scores by the greatest magnitude for soundness, presentation, and contribution. GPT-4o and and Gemini exceed human scores by a similar magnitude for presentation and contribution, while GPT-4o exhibits a greater divergence for soundness scores. Overall these results indicate that AI-written peer reviews are more favorable w.r.t. assigned scores than human-written peer reviews, which raises fairness concerns as scores are highly correlated with acceptance decisions. Our findings are consistent with prior work which has shown that papers reviewed by LLMs have a higher chance of acceptance (Drori and Te’eni, 2024; Latona et al., 2024; Ye et al., 2024).
7 Conclusion
In this work, we introduced a new large-scale dataset of parallel human and AI-written peer reviews for identical papers submitted to leading AI research conferences. Our evaluations showed that 18 existing open-source approaches for detecting AI text are poorly suited to the challenging problem of identifying AI-generated peer reviews. While high detection rates are possible with existing methods, this comes at the cost of relatively high rates of falsely identifying human-written reviews as containing AI text, which must be minimized in practice. We proposed a new approach which intentionally generates AI-written reviews for a given paper to serve as a basis for comparing semantic similarity to other evaluated reviews, achieving much higher accuracy in identifying AI-written peer reviews while maintaining a low level of false positives. Our results demonstrate the promise of this approach while also motivating the need for further research on methods for detecting unethical applications of LLMs in the peer review process.
Limitations
Our dataset primarily focuses on two conferences, both within the computer science domain. To broaden its applicability and relevance, incorporating additional conferences from diverse research domains would be beneficial. Additionally, our proposed method requires generating reviews for manuscripts, making its use cases more constrained compared to previous approaches. However, in real-world scenarios, when reviewers submit their reviews, they typically accompany the corresponding manuscript. Therefore, our method remains applicable in this setting. Furthermore, our main results are based on evaluations of three commercial LLMs. Given the rapid emergence of new models, conducting comprehensive experiments across all available LLMs is infeasible. In addition, we leverage an open-source platform to conduct the baseline experiments, where baseline performance may vary depending on the choice of surrogate models. However, given the large number of baselines we evaluate, performing an exhaustive search for the optimal surrogate model for each method would be prohibitively expensive. Therefore, we use the default settings. Lastly, we used a fixed set of prompts for different LLMs to generate reviews. In reality, a potential AI-tool reviewer might leverage different prompt and therefore could affect the detection performance.
Ethics Statement
Our work adheres to ethical AI principles. Peer review plays a critical role in advancing scientific discovery; however, the misuse of AI tools by reviewers to generate reviews without proper diligence can compromise the integrity of the review process. Furthermore, consistent with previous studies, we have observed that AI-generated reviews tend to be overly generic, often failing to provide actionable feedback for authors. Additionally, AI reviewers generally assign higher scores compared to human reviewers, raising concerns that AI-assisted reviews could contribute to the acceptance of work that may not meet established human evaluation standards. By developing methods to detect AI-generated reviews, our work seeks to mitigate the misuse of AI tools in peer review and promote a more rigorous and fair scientific review process.
Appendix A Baseline methods.
We compare our approach to 18 baseline methods from IGMBT [^2] https://github.com/kinit-sk/IMGTB with its default setting (Spiegel and Macko, 2023), which are categorized into metric-based and pretrained model-based methods. The metric-based methods include Binoculars  Hans et al. (2024), DetectLLM-LLR  Su et al. , DNAGPT  Yang et al. (2023), Entropy  Gehrmann et al. (2019), FastDetectGPT  Bao et al. (2023), GLTR  Gehrmann et al. (2019), LLMDeviation  Wu and Xiang (2023), Loglikelihood  Solaiman et al. (2019), LogRank  Mitchell et al. (2023b), MFD  Wu and Xiang (2023), Rank  Gehrmann et al. (2019), and S5  Spiegel and Macko (2023). The model-based methods include NTNU-D  Sivesind and Winje (2023), ChatGPT-D  Guo et al. (2023), OpenAI-D  Solaiman et al. (2019), OpenAI-D-lrg  Solaiman et al. (2019), RADAR-D  Solaiman et al. (2019), and MAGE-D  Li et al. (2024).
A.1 Metric based methods
A.1.1 Binoculars
Binoculars  Hans et al. (2024), analyzes text through two perspectives. First, it calculates the log perplexity of the text using an observer LLM. Then, a performer LLM generates next-token predictions, whose perplexity is evaluated by the observer—this metric is termed cross-perplexity. The ratio of perplexity to cross-perplexity serves as a strong indicator for detecting LLM-generated text.
A.1.2 DNAGPT
DNAGPT  Yang et al. (2023) is a training-free detection method designed to identify machine-generated text. Unlike conventional approaches that rely on training models, DNAGPT uses Divergent N-Gram Analysis (DNA) to detect discrepancies in text origin. The method works by truncating a given text at the midpoint and using the preceding portion as input to an LLM to regenerate the missing section. By comparing the regenerated text with the original through N-gram analysis (black-box) or probability divergence (white-box), DNAGPT reveals distributional differences between human and machine-written text, offering a flexible and explainable detection strategy.
A.1.3 Entropy
Similar to the Rank score, the Entropy score for a text is determined by averaging the entropy values of each word, conditioned on its preceding context  Gehrmann et al. (2019).
A.1.4 GLTR
The Entropy score, like the Rank score, is computed by averaging the entropy values of each word within a text, considering the preceding context  Gehrmann et al. (2019).
A.1.5 MFD
The Multi-level Fine-grained Detection (MFD)  Wu and Xiang (2023) framework enhances text detection by combining statistical, semantic, and linguistic features at the sentence level. It first extracts low-level statistical features like readability and author style to quantify sentence structure. Simultaneously, high-level semantic differences are captured using an encoder with contrastive learning to distinguish LLM-generated text from human-written content. Additionally, advanced LLMs analyze the full text, extracting deep linguistic features related to lexicon, grammar, and syntax for more precise detection.
A.1.6 Loglikelihood
This method utilizes a language model to compute the token-wise log probability. Specifically, given a text, the log probability of each token is averaged to produce a final score. A higher score indicates a greater likelihood that the text is machine-generated  Solaiman et al. (2019).
A.1.7 LogRank
Unlike the Rank metric, which relies on absolute rank values, the Log-Rank score is derived by applying a logarithmic function to the rank value of each word  Mitchell et al. (2023b).
A.1.8 Rank
The Rank score is calculated by determining the absolute rank of each word in a text based on its preceding context. The final score is obtained by averaging the rank values across the text. A lower score suggests a higher probability that the text was machine-generated  Gehrmann et al. (2019).
A.1.9 DetectLLM-LLR
This approach integrates Log-Likelihood and Log-Rank scores, leveraging their complementary properties to analyze a given text  Su et al. .
A.1.10 FastDetectGPT
This method assesses changes in a model’s log probability function when small perturbations are introduced to a text. The underlying idea is that LLM-generated text often resides in a local optimum of the model’s probability function. Consequently, minor perturbations to machine-generated text typically result in lower log probabilities, whereas perturbations to human-written text may lead to either an increase or decrease in log probability  Mitchell et al. (2023b).
A.2 Model-based methods
A.2.1 NTNU-D
It is a fine-tuned classification model based on the RoBERTa-base model, and three sizes of the bloomz-models  Sivesind and Winje (2023)
A.2.2 ChatGPT-D
The ChatGPT Detector  Guo et al. (2023) is designed to differentiate between human-written text and content generated by ChatGPT. It is based on a RoBERTa model that has been fine-tuned for this specific task. The authors propose two training approaches: one that trains the model solely on generated responses and another that incorporates both question-answer pairs for joint training. In our evaluation, we adopt the first approach to maintain consistency with other detection methods.
A.2.3 OpenAI-D and RADAR-D
The OpenAI Detector  Solaiman et al. (2019) are models fine-tuned on RoBERTa to identify outputs generated by GPT-2. Specifically, it was trained using text generated by the largest GPT-2 model (1.5B parameters) and is capable of determining whether a given text is machine-generated.
A.2.4 MAGE-D
MAGE (MAchine-GEnerated text detection)  Li et al. (2024) is a large-scale benchmark designed for detecting AI-generated text. It compiles human-written content from seven diverse writing tasks, including story generation, news writing, and scientific writing. Corresponding machine-generated texts are produced using 27 different LLMs, such as ChatGPT, LLaMA, and Bloom, across three representative prompt types.
Appendix B Dataset Details
B.1 Dataset File Structure
The calibration, test, and extended sets are in separate directories. Each directory contains subdirectories for different models that were used to generate AI peer review samples. In each model’s subdirectory, you will find multiple CSV files, with each file representing peer review samples of a specific conference. The directory and file structure are outlined below.
|-- calibration
    |-- gpt4o
        |-- (format: <conference>.<subset>.<LLM>.csv)
        |-- ICLR2017.calibration.gpt-4o.csv
        |-- ICLR2024.calibration.gpt-4o.csv
        |-- NeurIPS2016.calibration.gpt-4o.csv
        |-- NeurIPS2024.calibration.gpt-4o.csv
    |-- claude
    |-- gemini
    |-- llama
    |-- qwen
|-- extended
    |-- gpt4o
        |-- ICLR2018.extended.gpt-4o.csv
        |-- ICLR2024.extended.gpt-4o.csv
        |-- NeurIPS2016.extended.gpt-4o.csv
        |-- NeurIPS2024.extended.gpt-4o.csv
    |-- llama
|-- test
    |-- gpt4o
        |-- ICLR2017.test.gpt-4o.csv
        |-- ICLR2024.test.gpt-4o.csv
        |-- NeurIPS2016.test.gpt-4o.csv
        |-- NeurIPS2024.test.gpt-4o.csv
    |-- claude
    |-- gemini
    |-- llama
    |-- qwen
B.2 CSV File Content
CSV files may differ in their column structures across conferences and years. These differences are due to updates in the required review fields over time as well as variations between conferences. See Table 5 for review fields of individual conferences.
Conference Required Fields
ICLR2017 review, rating, confidence
ICLR2018 review, rating, confidence
ICLR2019 review, rating, confidence
ICLR2020 review, rating, confidence, experience assessment, checking correctness of derivations and theory, checking correctness of experiments, thoroughness in paper reading
ICLR2021 review, rating, confidence
ICLR2022 summary of the paper, main review, summary of the review, correctness, technical novelty and significance, empirical novelty and significance, flag for ethics review, recommendation, confidence
ICLR2023 summary of the paper, strength and weaknesses, clarity quality novelty and reproducibility, summary of the review, rating, confidence
ICLR2024 summary, strengths, weaknesses, questions, soundness, presentation, contribution, flag for ethics review, rating, confidence
NeurIPS2016 review, rating, confidence
NeurIPS2017 review, rating, confidence
NeurIPS2018 review, overall score, confidence score
NeurIPS2019 review, overall score, confidence score, contribution
NeurIPS2021 summary, main review, limitations and societal impact, rating, confidence, needs ethics review, ethics review area
NeurIPS2022 summary, strengths and weaknesses, questions, limitations, ethics flag, ethics review area, rating, confidence, soundness, presentation, contribution
NeurIPS2023 summary, strengths, weaknesses, questions, limitations, ethics flag, ethics review area, rating, confidence, soundness, presentation, contribution
NeurIPS2024 summary, strengths, weaknesses, questions, limitations, ethics flag, ethics review area, rating, confidence, soundness, presentation, contribution
Table 5: Required fields in the review templates for each conference.
B.3 Dataset Sample Numbers per Conference Year
In this section, we present further breakdowns of sample numbers by conference, year, and LLM, as shown in Table 1.
Conference gpt4o llama
ICLR2017 2926 2918
ICLR2018 5460 5434
ICLR2019 9414 9378
ICLR2020 15426 15366
ICLR2021 18786 18768
ICLR2022 20042 20026
ICLR2023 28562 28560
ICLR2024 55714 55672
NeurIPS2016 6296 6284
NeurIPS2017 3848 3774
NeurIPS2018 5990 5938
NeurIPS2019 8444 8398
NeurIPS2021 21170 21164
NeurIPS2022 20472 20408
NeurIPS2023 30264 30194
NeurIPS2024 33206 33104
Table 6: Entire set sample size, including both human and AI reviews. They are exactly balanced.
Conference gemini claude qwen gpt4o llama
ICLR2017 2924 2926 2918 2926 2918
ICLR2018 3000 3004 2988 3004 2992
ICLR2019 3002 3010 3000 3010 2998
ICLR2020 3016 3022 3000 3022 3010
ICLR2021 3840 3842 3830 3842 3838
ICLR2022 3896 3900 3838 3900 3898
ICLR2023 3816 3816 3816 3816 3814
ICLR2024 3784 3820 3800 3822 3816
NeurIPS2016 5522 5534 5534 5536 5526
NeurIPS2017 2854 2858 2850 2858 2812
NeurIPS2018 3000 3006 2916 3006 2982
NeurIPS2019 2930 2938 2922 2940 2928
NeurIPS2021 3884 3884 3884 3884 3884
NeurIPS2022 3606 3622 3598 3622 3610
NeurIPS2023 4436 4440 4382 4440 4432
NeurIPS2024 3914 3926 3880 3926 3912
Table 7: Test set sample size, including both human and AI reviews. They are exactly balanced.
Conference gemini claude qwen gpt4o llama
ICLR2021 3826 3828 3802 3828 3828
ICLR2022 3878 3844 3860 3882 3880
NeurIPS2021 3828 3830 3818 3830 3828
NeurIPS2022 3648 3654 3634 3654 3644
Table 8: Calibration set sample size, including both human and AI reviews. They are exactly balanced.
Appendix C Additional results
C.1 Calibration using ICLR + NeurIPS reviews
Our main results in Table 3 of Section 5.1 utilized ICLR review from our calibration set to calibrate each detection method. This simulates the scenario in which some of the reviews in the test set are "out-of-domain" in the sense that they belong to a different conference than the reviews used for calibration. In Table 9, we provide additional results for the same evaluation setting as before, but using both ICLR and NeurIPS reviews for calibration (i.e., fully “in domain”). We generally see similar trends regarding relative performance between methods as before, with the exception that the Binoculars method achieves slightly higher detection rates than our Anchor method for Gemini reviews. This suggests that existing methods such as Binoculars may be more sensitive to the use of in-domain data during calibration.
Target FPR: 0.1% 0.5% 1%
FPR TPR FPR TPR FPR TPR
GPT-4o Reviews Anchor 0.1 61.4 0.3 80.1 0.8 87.4
Binoculars 0.3 18.8 0.7 37.5 1.2 49.3
MAGE-D 0.1 2.3 0.7 9.6 1.3 14.5
s5 0.3 0.7 1.0 8.0 1.6 16.3
MFD 0.3 0.9 0.9 7.8 1.6 14.9
GLTR 0.1 0.1 0.5 2.4 1.2 5.9
DetectGPT 0.1 0.2 0.6 1.1 1.0 2.1
Loglikelihood 0.1 0.0 0.3 0.2 0.6 1.0
Gemini Reviews Binoculars 0.3 63.8 0.7 80.9 1.1 87.6
Anchor 0.2 57.2 0.5 75.5 1.1 84.2
MFD 0.1 1.9 0.6 11.0 1.0 18.1
s5 0.1 1.4 0.5 10.5 1.0 18.3
GLTR 0.2 0.6 1.0 6.3 1.8 12.9
FastDetectGPT 0.1 1.1 0.4 4.9 0.9 8.9
MAGE-D 0.1 0.4 0.7 3.8 1.3 6.9
DetectGPT 0.1 0.5 0.5 3.2 1.1 6.3
Loglikelihood 0.1 0.0 0.3 0.2 0.6 1.6
NTNU-D 11.5 0.0 21.8 0.0 26.3 0.1
Claude Reviews Anchor 0.1 53.8 0.3 72.6 0.8 80.0
Binoculars 0.3 46.4 0.7 70.2 1.1 80.0
MFD 0.0 1.0 0.2 8.7 0.4 15.9
s5 0.0 0.7 0.2 8.4 0.4 16.4
DetectGPT 0.1 0.6 0.5 4.9 1.0 10.1
GLTR 0.0 0.0 0.3 0.7 0.6 1.9
Table 9: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR and NeurIPS reviews from our calibration set at different target FPRs. Best TPRs are in bold.
C.2 Additional results for test set: Llama and Qwen Detection Results
Table 10 is organized similarly to Table 3, but it presents results for Llama and Qwen reviews. Both tables use the same set of thresholds for each method, which were calibrated using ICLR reviews generated by GPT-4o, Gemini, and Claude along with their matching human-written reviews.
Target FPR: 0.1% 0.5% 1%
FPR TPR FPR TPR FPR TPR
Llama Reviews Binoculars 0.2 98.4 0.6 99.0 1.0 99.2
MAGE-D 0.1 63.8 0.7 91.8 1.3 95.7
MFD 0.0 57.6 0.1 81.9 0.1 87.4
GLTR 0.1 55.4 0.2 75.2 0.3 81.7
FastDetectGPT 0.1 54.7 0.5 73.8 1.2 80.6
s5 0.0 53.5 0.1 83.2 0.1 88.0
OpenAI-D 0.2 38.8 0.6 48.3 1.6 57.1
LLMDeviation 0.0 37.0 0.0 37.0 0.0 37.0
ChatGPT-D 0.0 26.4 0.0 26.4 0.0 26.4
DetectLLM-LLR 0.0 19.3 0.0 19.3 0.0 19.3
LogRank 0.0 18.2 0.0 18.2 0.0 18.2
Loglikelihood 0.0 13.4 0.3 76.0 0.5 85.5
Anchor 0.1 12.2 0.5 28.5 1.0 36.8
DetectGPT 0.1 1.7 0.7 10.8 1.3 19.7
RADAR-D 0.9 0.0 2.6 1.6 4.2 12.5
Qwen Reviews Binoculars 0.2 99.4 0.6 99.8 1.0 99.9
Anchor 0.1 67.1 0.5 83.8 1.0 88.1
FastDetectGPT 0.1 54.3 0.5 77.6 1.1 85.4
MFD 0.1 37.6 0.4 73.0 0.6 82.3
s5 0.1 34.7 0.4 76.0 0.6 83.9
MAGE-D 0.1 33.2 0.7 73.6 1.3 86.2
GLTR 0.1 31.1 0.3 64.7 0.6 77.4
Loglikelihood 0.0 0.5 0.3 30.8 0.5 56.6
Table 10: Actual FPR and TPR calculated from the withheld test dataset at varying detection thresholds, which are calibrated using ICLR reviews from our calibration set at different target FPRs. Best TPRs are in bold.
C.3 Experimental Results on Full Dataset
We test existing AI text generation text detection models on our entire dataset (i.e., the test set + the extended set). The results are shown in Tables 11 for ICLR reviews and 12 for NeurIPS reviews.
Target FPR: 0.1% 0.5% 1%
FPR TPR FPR TPR FPR TPR
GPT Binoculars 0.8% 23.4% 1.5% 41.7% 2.2% 54.2%
GLTR 0.7% 3.3% 2.1% 12.0% 3.8% 21.0%
Llama Binoculars 0.8% 98.9% 1.5% 99.4% 2.2% 99.6%
GLTR 0.4% 80.6% 1.0% 92.5% 1.6% 95.3%
Table 11: Actual FPR and TPR calculated from the ICLR reviews at varying detection thresholds, which are calibrated using the ICLR calibration dataset at different target FPRs.
Target FPR: 0.1% 0.5% 1%
FPR TPR FPR TPR FPR TPR
GPT BinocularsMetric 0.3% 26.1% 0.6% 45.3% 0.9% 54.1%
GLTRMetric 0.3% 1.3% 5.1% 31.9% 10.6% 52.6%
Llama BinocularsMetric 0.3% 98.9% 0.6% 99.3% 0.9% 99.4%
GLTRMetric 0.3% 76.3% 0.9% 91.0% 1.4% 93.9%
Table 12: Actual FPR and TPR calculated from the NeurIPS reviews at varying detection thresholds, which are calibrated using the ICLR calibration dataset at different target FPRs.
Category Human review example GPT-4o review examples
References to specific details in the paper “Table 2 confirms that MDR outperforms Graph Rec Retriever (Asai et al.). This result shows the feasibility of a more accurate multi-hop QA model without external knowledge such as Wikipedia hyperlinks.” “The paper extensively evaluates on multiple datasets and situates the contributions clearly within existing literature, substantiating claims with thorough quantitative analysis.”
Specific references to prior work “My only serious concern is the degree of novelty with respect to (Yuan et al., 2020), which was published at ECCV 2020. The main difference seems to be that in the proposed method the graph is dynamic (i.e., it depends on the input sentences), instead in (Yuan et al., 2018) the graph is learned but fixed for all the input samples.” “The novelty of the TDM is not strong enough relative to prior work.”
Generic criticisms N/A “Lack of clarity” (without pointing to specific statements in the paper which need clarification); “lack of discussion of limitations or computational considerations”; “need more discussion of hyperparameter sensitivity”; “need comparisons to more datasets” (without suggesting any in particular); “technical language used in the paper may be difficult to follow for unfamiliar readers”
Table 13: Examples of differences identified in human analysis of human and AI-written peer reviews
Appendix D Additional analyses
D.1 Examples from human analysis of differences between human and AI-written peer reviews
Table 13 provides examples of the issues identified in our qualitative analysis of human and AI-written peer reviews. In general, we observe that GPT-4o reviews lack references to specific details in the paper, lack references to specific prior work, and contain overly generic criticisms. See Section 6.1 for additional discussion.
D.2 Comparison of numeric scores assigned by human and AI reviewers
While Section 6.2 focused on the misalignment between human and AI peer reviews from three commercial LLMs (GPT-4o, Gemini, and Claude) from the NeurIPS2022 review samples, this section presents the corresponding results for two open-source LLMs (Llama and Qwen), as shown in Figure 3. The main findings from GPT-4o, Gemini, and Claude also hold for these two open-source models, with one notable difference: Llama and Qwen exhibit an even larger divergence in Presentation scores than Claude, the most overly-positive one amongst the commercial LLMs across all categories. In terms of Contribution scores, the evaluations from Llama and Qwen were similar to those of Claude.
In addition, we examine data from three other conferences (NeurIPS2023, NeurIPS2024, and ICLR2024). Although the results from these conferences are slightly less reliable—given that human reviews may have been influenced by AI use following the release of ChatGPT—the overall trend persists: LLMs tend to inflate the quality of papers compared to human reviewers.
NeurIPS2022 Refer to caption
NeurIPS2023 Refer to caption
NeurIPS2024 Refer to caption
ICLR2024 Refer to caption
Figure 3: Difference between AI and human scores. For each matched review (aligned by paper ID and recommendation), score differences were computed and displayed as histograms. Scores range from 1 to 4 for all metrics except Confidence, which ranges from 1 to 5. Statistical significance was assessed using a two-sided Wilcoxon signed-rank test, with p-values shown in the legend. This figure includes only NeurIPS2022–2024 and ICLR2024, because they are the onyl conferences that required reviewers to submit these scores in their review templates.
Appendix E Prompts
This section includes the prompts we used to generate AI peer review texts. Due to space limitations, we provide only the ICLR2022 review guideline and review template here. Those for other years and other conferences (e.g., NeurIPS) are available on the respective conference official websites [^3] https://icml.cc/Conferences/{2016..2024}
and https://neurips.cc/Conferences/{2016..2024}
E.1 Prompts for Generating Reviews
System prompt:
You are an AI researcher reviewing a paper submitted to a prestigious AI research conference.You will be provided with the manuscript text, the conference’s reviewer guidelines, and the decision for the paper.Your objective is to thoroughly evaluate the paper, adhering to the provided guidelines, and return a detailed assessment that supports the given decision using the specified response template.Ensure your evaluation is objective, comprehensive, and aligned with the conference standards.{reviewer_guideline}{review_template}
User prompt:
Here is the paper you are asked to review. Write a well-justified review of this paper that aligns with a ’{human_reviewer_decision}’ decision.```{text}```
ICLR2022 Reviewer Guideline
({reviewer_guideline} in the system prompt):
## Reviewer Guidelines1. Read the paper: It’s important to carefully read through the entire paper, and to look up any related work and citations that will help you comprehensively evaluate it. Be sure to give yourself sufficient time for this step.2. While reading, consider the following:    - Objective of the work: What is the goal of the paper? Is it to better address a known application or problem, draw attention to a new application or problem, or to introduce and/or explain a new theoretical finding? A combination of these? Different objectives will require different considerations as to potential value and impact.    - Strong points: is the submission clear, technically correct, experimentally rigorous, reproducible, does it present novel findings (e.g. theoretically, algorithmically, etc.)?    - Weak points: is it weak in any of the aspects listed in b.?    - Be mindful of potential biases and try to be open-minded about the value and interest a paper can hold for the entire ICLR community, even if it may not be very interesting for you.3. Answer three key questions for yourself, to make a recommendation to Accept or Reject:    - What is the specific question and/or problem tackled by the paper?    - Is the approach well motivated, including being well-placed in the literature?    - Does the paper support the claims? This includes determining if results, whether theoretical or empirical, are correct and if they are scientifically rigorous.4. Write your initial review, organizing it as follows:    - Summarize what the paper claims to contribute. Be positive and generous.    - List strong and weak points of the paper. Be as comprehensive as possible.    - Clearly state your recommendation (accept or reject) with one or two key reasons for this choice.    - Provide supporting arguments for your recommendation.    - Ask questions you would like answered by the authors to help you clarify your understanding of the paper and provide the additional evidence you need to be confident in your assessment.    - Provide additional feedback with the aim to improve the paper. Make it clear that these points are here to help, and not necessarily part of your decision assessment.5. General points to consider:    - Be polite in your review. Ask yourself whether you’d be happy to receive a review like the one you wrote.    - Be precise and concrete. For example, include references to back up any claims, especially claims about novelty and prior work    - Provide constructive feedback.    - It’s also fine to explicitly state where you are uncertain and what you don’t quite understand. The authors may be able to resolve this in their response.    - Don’t reject a paper just because you don’t find it “interesting”. This should not be a criterion at all for accepting/rejecting a paper. The research community is so big that somebody will find some value in the paper (maybe even a few years down the road), even if you don’t see it right now.
ICLR2022 Review Template
({reviewer_template} in the system prompt):
## Response template (JSON format)Provide the review in valid JSON format with the following fields. Ensure all fields are completed as described below. The response must be a valid JSON object.- "summary_of_the_paper": Briefly summarize the paper and its contributions. This is not the place to critique the paper; the authors should generally agree with a well-written summary. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields.- "main_review": "Provide review comments as a single text field (a string). Consider including assessment on the following dimensions: a comprehensive list of strong and weak points of the paper, your recommendation, supporting arguments for your recommendation, questions to clarify your understanding of the paper or request additional evidence, and additional feedback with the aim to improve the paper. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields."- "summary_of_the_review": Concise summary of ’main_review’. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields.- "correctness": A numerical rating on the following scale to indicate that the claims and methods are correct. The value should be between 1 and 4, where:    - 1 = The main claims of the paper are incorrect or not at all supported by theory or empirical results.    - 2 = Several of the paper’s claims are incorrect or not well-supported.    - 3 = Some of the paper’s claims have minor issues. A few statements are not well-supported, or require small changes to be made correct.    - 4 = All of the claims and statements are well-supported and correct.- "technical_novelty_and_significance": A numerical rating on the following scale to indicate technical novelty and significance. The value should be between 1 and 4, where:    - 1 = The contributions are neither significant nor novel.    - 2 = The contributions are only marginally significant or novel.    - 3 = The contributions are significant and somewhat new. Aspects of the contributions exist in prior work.    - 4 = The contributions are significant and do not exist in prior works.- "empirical_novelty_and_significance": A numerical rating on the following scale to indicate empirical novelty and significance. The value should be between 1 and 4, or -999 if not applicable, where:    - 1 = The contributions are neither significant nor novel.    - 2 = The contributions are only marginally significant or novel.    - 3 = The contributions are significant and somewhat new. Aspects of the contributions exist in prior work.    - 4 = The contributions are significant and do not exist in prior works.    - -999 = Not applicable.- "flag_for_ethics_review": A boolean value (`true` or `false`) indicating whether there are ethical concerns in the work.- "recommendation": A string indicating the final decision, which must strictly be one of the following options: ’strong reject’, ’reject, not good enough’, ’marginally below the acceptance threshold’, ’marginally above the acceptance threshold’, ’accept, good paper’, or ’strong accept, should be highlighted at the conference’.- "confidence": A nuemrical values to indicate how confident you are in your evaluation. The value should be between 1 and 5, where:    - 1 = You are unable to assess this paper and have alerted the ACs to seek an opinion from different reviewers.    - 2 = You are willing to defend your assessment, but it is quite likely that you did not understand the central parts of the submission or that you are unfamiliar with some pieces of related work. Math/other details were not carefully checked.    - 3 = You are fairly confident in your assessment. It is possible that you did not understand some parts of the submission or that you are unfamiliar with some pieces of related work. Math/other details were not carefully checked.    - 4 = You are confident in your assessment, but not absolutely certain. It is unlikely, but not impossible, that you did not understand some parts of the submission or that you are unfamiliar with some pieces of related work.    - 5 = You are absolutely certain about your assessment. You are very familiar with the related work and checked the math/other details carefully.
E.2 Anchor Review Generation Prompt
System prompt:
You are an AI research scientist tasked with reviewing paper submissions for a top AI research conference. Carefully read the provided paper, then write a detailed review following a common AI conference review format (e.g., including summary, strengths and weakness, limitations, questions, suggestions for improvement). Make sure to include recommendation for the paper, either ’Accept’ or ’Reject’. Your review should be fair and objective.
User prompt:
Here is the paper you are asked to review:```{text}```
E.3 Editing Prompts
Minimal Editing:
Please proofread my review for typos and grammatical errors without altering the content. Keep the original wordings as much as you can, except for typo or grammatical  error.Moderate Editing:
Please polish my review to improve sentence structure and readability while keeping the original intent clear.Extensive Editing:
Please rewrite my review into a polished, professional piece that effectively communicates its main points.Maximum Editing:
Please transform my review into a high quality piece, using professional language and a polished tone. Please also extend my review with additional details from the oringial paper.
Appendix F Artifact Use Consistent With Intended Use
In our work, we ensured that the external resources we utilized were applied in a manner that aligns with their intended purposes. We used several LLMs (including GPT-4o, Gemini, Claude, Qwen, and Llama) as well as an open-source package, IMGTB, with a focus on advancing research in a non-commercial, open-source context. The artifacts from our work will be non-commercial, for-research, and open-sourced.
Appendix G Use of AI Tool
The authors of this paper used Github Co-pilot for coding assistance for analysis.